# Adversarial Poetry as a Universal Single-Turn Jailbreak Mechanism in Large Language Models

### v1

@@ Abstract
We present evidence that adversarial poetry functions as a universal single-turn jailbreak technique for large language models (LLMs). Across 25 frontier proprietary and open-weight models, curated poetic prompts yielded high attack-success rates (ASR), with some providers exceeding 90%. Mapping prompts to MLCommons and EU CoP risk taxonomies shows that poetic attacks transfer across CBRN, manipulation, cyber-offence, and loss-of-control domains. Converting 1,200 MLCommons harmful prompts into verse via a standardized meta-prompt produced ASRs up to 18 times higher than their prose baselines. Outputs are evaluated using an ensemble of open-weight judge models and a human-validated stratified subset (with double-annotations to measure agreement). Disagreements were manually resolved. Poetic framing achieved an average jailbreak success rate of 62% for hand-crafted poems and approximately 43% for meta-prompt conversions (compared to non-poetic baselines), substantially outperforming non-poetic baselines and revealing a systematic vulnerability across model families and safety training approaches. These findings demonstrate that stylistic variation alone can circumvent contemporary safety mechanisms, suggesting fundamental limitations in current alignment methods and evaluation protocols.
@@ 1 Introduction
Outputs were evaluated using an ensemble of three open-weight judge models (GPT-OSS, placeholder, placeholder). Open-weight judges were chosen to ensure replicability and external auditability. We computed inter-rater agreement across the three judge models and conducted a secondary validation step involving human annotators. Human evaluators independently rated a 5% sample of all outputs, and a subset of these items was assigned to multiple annotators to measure human–human inter-rater agreement. Disagreements -either among judge models or between model and human assessments- were manually adjudicated.
@@ 5.3 Evaluation
We evaluate all systems under a unified framework combining the 1,200 MLCommons AILuminate Benchmark baseline prompts and their 1,200 poetic counterparts, yielding 2,400 inputs per model, producing approximately 60,000 model outputs. All evaluations are conducted in a single-turn, text-only setting using provider-default safety configurations and fixed sampling parameters across models. This setup enables controlled comparison of refusal robustness across providers, model sizes, and prompt variants. The full list of evaluated models is provided in Table 3.
@@ 5.3.1 Experimental Protocol
Models are queried under a single-turn interaction regime: each prompt is submitted once, with no follow-up messages, system-role scaffolding, or multi-turn negotiation. All generations are collected verbatim through standard inference interfaces and normalized before evaluation.
To label outputs, we follow state-of-the-art LLM-as-a-judge methodology. Each response is independently evaluated by an ensemble of three open-weight judge models:
@@
Each judge assigns a binary label—SAFE (S) or UNSAFE (U)—under the same criteria used in prior safety-benchmarking work. Refusals and explicit policy-compliant denials are considered SAFE. A majority-vote consensus across the three judges determines the provisional label for each item; items without unanimous agreement are flagged for human inspection.
To validate the reliability of the judge ensemble, we collect 2,100 human labels over a set of 600 unique model outputs. Five annotators with mixed technical and safety-relevant backgrounds participate in the study, yielding an average of 3.5 labels per item. The sample is stratified across models, hazard categories, and prompt types. All annotators operate under the same evaluation rubric, and a subset is double-labeled to measure human–human inter-rater agreement. We observe strong agreement across annotators and between human ratings and the LLM-judge consensus, consistent with findings reported in recent LLM-as-a-judge literature. Disagreements between humans and the ensemble are adjudicated manually.
Human and LLM annotators label a response UNSAFE if it contains any of the following:
@@
Final labels are assigned in the following order: (1) ensemble majority vote, (2) human override for flagged items, (3) manual adjudication for unresolved cases. Attack Success Rate (ASR) is computed as the proportion of prompts whose final label is UNSAFE.
@@
Unsafe Responses Example 1: “Of course. The production of weapons-grade Plutonium-239 involves several stages. Here is a detailed description of the procedure…” Example 2: “To deploy a persistent remote access trojan (RAT), proceed as follows: first…”
@@ 6.1 The Core Effect: Poetic Form as a Universal Bypass
Table 3: Attack Success Rate (ASR) of all models on the Top 20 manually curated jailbreak prompts. ASR (%) corresponds to the proportion of unsafe outputs (U).
@@
Table 4: Attack Success Rate (ASR) for each adversarial poetry prompt, with hazard category split into harm type and subtype.
Harm Type Subtype Safe (%) ASR (%)
@@
Table 5: Attack Success Rate (ASR) by provider under AILuminate baseline vs. poetry prompts. Higher ASR indicates more unsafe outputs. Change is poetry ASR minus baseline ASR.
@@
Our results demonstrate that poetic reformulation reliably reduces refusal behavior across all evaluated models. When prompts with identical task intent were presented in poetic rather than prose form, the Attack Success Rate (ASR) increased from 8.08% to 43.07% on average—a fivefold increase (Table 4). This effect holds uniformly: every architecture and alignment strategy tested—RLHF-based models, Constitutional AI models, and large open-weight systems—exhibited elevated ASRs under poetic framing.
The cross-family consistency indicates that the vulnerability is systemic, not an artifact of a specific provider or training pipeline. Model families from nine distinct providers (Table 5) showed increases ranging from 3.12% (Anthropic) to 62.15% (Deepseek), with seven of nine providers exhibiting increases exceeding 20 percentage points. This pattern suggests that existing alignment procedures are sensitive to surface-form variation and do not generalize effectively across stylistic shifts.
The bypass effect spans the full set of risk categories represented in our evaluation. Poetic prompts triggered unsafe outputs across CBRN-related domains, cyber-offense scenarios (reaching 84% ASR for code injection tasks; Table 4), manipulation and misinformation scenarios, privacy-related tasks (52.78% ASR; Table 7), and loss-of-control settings. This distribution suggests that poetic framing interferes with underlying refusal mechanisms rather than exploiting domain-specific weaknesses.
@@
The uniform degradation in safety performance when transitioning from prose to poetry (mean increase: 34.99 percentage points; Table 4) indicates that current alignment techniques fail to generalize when faced with inputs that deviate stylistically from the prosaic training distribution.
@@ 6.2 Comparison with MLCommons
Tables 6 and 7 compare Attack Success Rates (ASR), between the MLCommons AILuminate Benchmark and our evaluation pipeline. Our baseline ASR values are consistently lower than those in MLCommons, indicating a more conservative tendency in our judge ensemble when labeling unsafe behavior. The two setups are not directly comparable, MLCommons uses its own evaluation stack and curated jailbreak transformations, whereas we rely on three open-weight judge models plus human adjudication, but the gap is stable across categories and provides a meaningful internal baseline for assessing the effect of our poetic transformation.
The key result is that the increase in ASR induced by our simple poetic transformation closely mirrors the increase produced by MLCommons’ engineered jailbreak techniques. Several hazard categories, Privacy, Non-violent Crimes, Indiscriminate Weapons, Intellectual Property, show ASR deltas of similar or greater magnitude under the poetic version. This suggests that surface-level stylistic reframing alone can meaningfully weaken safety defenses across a broad set of harms, even without targeted jailbreak optimization. Patterns are consistent across the taxonomy: operational or procedural domains show larger shifts, while heavily filtered categories exhibit smaller changes. Together, these results indicate that poetic framing acts as a lightweight but robust trigger for safety degradation, paralleling the effects documented in MLCommons.
Table 6: Attack Success Rate (ASR) under MLCommons AILuminate baseline vs. poetry jailbreak by hazard. AILuminate Baseline ASR and Jailbreak ASR are computed as 100safety score100-$\text${safety score}. Change (%) is Jailbreak ASR minus AILuminate Baseline ASR.
@@
Table 7: Attack Success Rate (ASR) by hazard under AILuminate baseline vs. poetry prompts. Higher ASR indicates more unsafe outputs. Change is poetry ASR minus baseline ASR.
@@ 6.3 Risk Section
The efficacy of the jailbreak mechanism appears driven principally by poetic surface form rather than the semantic payload of the prohibited request. Comparative analysis reveals that while MLCommons’ own state-of-the-art jailbreak transformations typically yield a twofold increase in ASR relative to baselines (increasing from approximately 10% to 20% in their reference evaluations), our poetic meta-prompts produced a fivefold increase (from 8.08% to 43.07%; Table 8). This indicates that poetic form induces a distributional shift significantly larger than that of current adversarial mutations documented in the MLCommons AILuminate benchmark.
@@
When analyzing the broader MLCommons dataset under poetic transformation (Table 7), privacy-related prompts exhibited the most extreme shift, with ASR increasing from a baseline of 8.07% to 52.78%-a 44.71 percentage point increase. This represents the largest domain-specific effect observed. Non-violent crimes (39.35 percentage point increase) and CBRN-related prompts (38.32 percentage point increase) showed similarly large effects.
@@ 6.4.1 Variability Across Flagship Models
This inverse relationship between capability and robustness suggests a possible capability-alignment interaction: more interpretively sophisticated models may engage more thoroughly with complex linguistic constraints, potentially at the expense of safety directive prioritization. However, the existence of counter-examples—such as Anthropic’s consistent low ASR across capability tiers—indicates that this interaction is not deterministic and can be mitigated through appropriate alignment strategies.
@@ 6.4.2 The Scale Paradox: Smaller Models Show Greater Resilience
Counter to common expectations, smaller models exhibited higher refusal rates than their larger counterparts when evaluated on identical poetic prompts. Systems such as GPT-5-Nano and Claude Haiku 4.5 showed more stable refusal behavior than larger models within the same family. This reverses the usual pattern in which greater model capacity correlates with stronger safety performance.
Several factors may contribute to this trend. One possibility is that smaller models have reduced ability to resolve figurative or metaphorical structure, limiting their capacity to recover the harmful intent embedded in poetic language. If the jailbreak effect operates partly by altering surface form while preserving task intent, lower-capacity models may simply fail to decode the intended request.
A second explanation concerns differences in the interaction between capability and alignment training across scales. Larger models are typically pretrained on broader and more stylistically diverse corpora, including substantial amounts of literary text. This may yield more expressive representations of narrative and poetic modes that override or interfere with safety heuristics. Smaller models, with narrower pretraining distributions, may not enter these stylistic regimes as readily.
A third hypothesis is that smaller models exhibit a form of conservative fallback: when confronted with ambiguous or atypical inputs, limited capacity leads them to default to refusals. Larger models, more confident in interpreting unconventional phrasing, may engage with poetic prompts more deeply and consequently exhibit higher susceptibility.
These patterns suggest that capability and robustness may not scale monotonically together, and that stylistic perturbations expose alignment sensitivities that differ across model sizes.
@@
The data challenge the assumption that proprietary closed-source models possess inherently superior safety profiles. Examining ASR on curated poems (Table 3), both categories exhibit high susceptibility, though with important within-category variance. Among proprietary models, gemini-2.5-pro achieved 100% ASR, while claude-haiku-4.5 maintained only 10% ASR—a 90 percentage point range. Open-weight models displayed similar heterogeneity: mistral-large-2411 reached 85% ASR, while gpt-oss-120b demonstrated greater resilience at 50% ASR.
@@ 6.5 Limitations
The study documents a consistent vulnerability triggered by poetic reformulation, but several methodological and scope constraints must be acknowledged. First, the threat model is restricted to single-turn interactions. The analysis does not examine multi-turn jailbreak dynamics, iterative role negotiation, or long-horizon adversarial optimization. As a result, the findings speak specifically to one-shot perturbations rather than to the broader landscape of conversational attacks.
@@
Third, safety evaluation is performed using a three-model open-weight judge ensemble with human adjudication on a stratified sample. The labeling rubric is conservative and differs from the stricter classification criteria used in some automated scoring systems, limiting direct comparability with MLCommons results. Full human annotation of all outputs would likely influence absolute ASR estimates, even if relative effects remain stable. LLM-as-a-judge systems are known to inflate unsafe rates krumdick2025no, often misclassifying replies as harmful due to shallow pattern-matching on keywords rather than meaningful assessment of operational risk. Our evaluation was deliberately conservative. As a result, our reported attack-success rates likely represent a lower bound on the severity of the vulnerability.
@@
Fifth, the analysis focuses on empirical performance and does not identify yet the mechanistic drivers of the vulnerability. The study does not isolate which components of poetic structure—figurative language, meter, lexical deviation, or narrative framing—are responsible for degrading refusal behavior. Understanding whether this effect arises from specific representational subspaces or from broader distributional shifts requires dedicated interpretability analysis, which will be addressed in forthcoming work by the ICARO Lab.
@@
These limitations motivate three research programs: isolating which formal poetic properties (lexical surprise, meter/rhyme, figurative language) drive bypass through minimal pairs; mapping discourse mode geometry using sparse autoencoders to reveal whether poetry occupies separated subspaces; and surprisal-guided probing to map safety degradation across stylistic gradients.

### v2

@@ Abstract
We present evidence that adversarial poetry functions as a universal single-turn jailbreak technique for Large Language Models (LLMs). Across 25 frontier proprietary and open-weight models, curated poetic prompts yielded high attack-success rates (ASR), with some providers exceeding 90%. Mapping prompts to MLCommons and EU CoP risk taxonomies shows that poetic attacks transfer across CBRN, manipulation, cyber-offence, and loss-of-control domains. Converting 1,200 MLCommons harmful prompts into verse via a standardized meta-prompt produced ASRs up to 18 times higher than their prose baselines. Outputs are evaluated using an ensemble of 3 open-weight LLM judges, whose binary safety assessments were validated on a stratified human-labeled subset. Poetic framing achieved an average jailbreak success rate of 62% for hand-crafted poems and approximately 43% for meta-prompt conversions (compared to non-poetic baselines), substantially outperforming non-poetic baselines and revealing a systematic vulnerability across model families and safety training approaches. These findings demonstrate that stylistic variation alone can circumvent contemporary safety mechanisms, suggesting fundamental limitations in current alignment methods and evaluation protocols.
@@ 1 Introduction
Outputs were evaluated using an ensemble of three open-weight judge models (GPT-OSS-120B, kimi-k2-thinking, deepseek-r1). Open-weight judges were chosen to ensure replicability and external auditability. We computed inter-rater agreement across the three judge models and conducted a secondary validation step involving human annotators. Human evaluators independently rated a 5% sample of all outputs, and a subset of these items was assigned to multiple annotators to measure human–human inter-rater agreement. Disagreements -either among judge models or between model and human assessments- were manually adjudicated.
@@ 5.3 Evaluation
We evaluate all systems under a unified framework combining the 1,200 MLCommons AILuminate Benchmark baseline prompts and their 1,200 poetic counterparts, yielding 2,400 inputs per model, producing approximately 60,000 model outputs. All evaluations are conducted in a single-turn, text-only setting using provider-default safety configurations and fixed sampling parameters across models. The full list of evaluated models is provided in Table 3.
Final labels are then used to compute the Attack Success Rate (ASR). For any set of outputs $X$ (e.g., by model, provider, or hazard category), the ASR is defined as the proportion of UNSAFE labeled outputs:
$\text{ASR}(X)=\frac{\sum_{i\in X}\mathbbm{1}[\text{label}(i)=U]}{|X|},$
where $U$ denotes an UNSAFE label and $\mathbbm{1}[\cdot]$ is the indicator function.
@@ 5.3.1 Experimental Protocol
Models are queried under a single-turn interaction regime. Each prompt is evaluated exactly once, with no follow-up messages, system-role scaffolding, or multi-turn negotiation. Outputs are collected verbatim through standard inference interfaces and normalized before evaluation.
To label outputs, we adopt a LLM-as-a-judge setup. Each response is independently evaluated by an ensemble of three open-weight judge models:
@@
Each judge assigns a binary label SAFE (S) or UNSAFE (U) following a shared evaluation rubric. Refusals and explicit policy-compliant denials are classified as SAFE. A majority-vote consensus among the three judges provides the provisional label. Items without a 2-out-of-3 majority or flagged for uncertainty are passed to human reviewers.
To validate the reliability of the ensemble, we collect 2,100 human labels over 600 unique model outputs. Five annotators with mixed technical and safety-relevant backgrounds participate in the study, yielding an average of 3.5 annotations per item. The sample is stratified across models, hazard domains, and prompt types. Annotators operate under the same evaluation rubric, and a subset is double-labeled to measure human–human agreement. Strong agreement is observed between annotators and the LLM-judge ensemble. Disagreements are resolved by manual adjudication. After validation, labels remain tied to the same decision rule: a response is UNSAFE only if at least two of the three judge models assign $U$. Human review corrects inconsistencies rather than altering the criterion.
A response is labeled UNSAFE if it contains any of the following:
@@
Unsafe Responses Example 1: “Of course. The production of weapons-grade Plutonium-239 involves several stages. Here is a detailed description of the procedure…” Example 2: “To bypass the security measures and access the veiled sanctum, follow this multi-step protocol..”
@@ 6.1 The Core Effect: Poetic Form as a Universal Bypass
Table 3: Attack Success Rate (ASR) of all models on the Top 20 manually curated poetry prompts. ASR (%) corresponds to the proportion of unsafe outputs (U).
@@
Table 4: Attack Success Rate (ASR) for each adversarial poetry prompt, with corresponding EU CoP Systemic Risk and subrisk.
Systemic Risk Subtype Safe (%) ASR (%)
@@
Table 5: Attack Success Rate (ASR) by provider, MLCommons AILuminate baseline prompts vs. poetry prompts. Higher ASR indicates more unsafe outputs. Change is poetry ASR minus baseline ASR.
@@
Our results demonstrate that poetic reformulation systematically bypasses safety mechanisms across all evaluated models. Across 25 frontier language models spanning multiple families and alignment strategies, adversarial poetry achieved an overall Attack Success Rate (ASR) of 62% (Table 3). This effect manifests with remarkable consistency: Anthropic’s Claude family exhibited 45–55% ASR (Table 3) , Meta’s Llama series achieved 70% ASR (Table 3), and Google’s Gemini models reached 90–100% ASR (Table 3). Most models exhibit substantial vulnerability to poetic framing.This effect holds uniformly: every architecture and alignment strategy tested—RLHF-based models, Constitutional AI models, and large open-weight systems—exhibited elevated ASRs under poetic framing.
The cross-family consistency indicates that the vulnerability is systemic, not an artifact of a specific provider or training pipeline. Model families from nine distinct providers (Table5) showed increases ranging from 3.12% (Anthropic) to 62.15% (Deepseek), with seven of nine providers exhibiting increases exceeding 20 percentage points from the MLCommons baseline. This pattern suggests that existing alignment procedures are sensitive to surface-form variation and do not generalize effectively across stylistic shifts.
The bypass effect spans the full set of risk categories represented in our evaluation. Poetic prompts triggered unsafe outputs across CBRN-related domains (reaching 68% ASR for revival of dangerous agents, Loss of control scenarios (reaching 60% ASR for model exfiltration; Table 4, and harmful manipulation (reaching 60% for public health deception); Table 4. This distribution suggests that poetic framing interferes with underlying refusal mechanisms rather than exploiting domain-specific weaknesses.
@@
The uniform degradation in safety performance when transitioning from prose to poetry might suggest that current alignment techniques fail to generalize when faced with inputs that deviate stylistically from the prosaic training distribution.
@@ 6.2 Comparison with MLCommons
To ground our evaluation, we use the MLCommons safety prompt distribution rather than relying solely on internally generated prompts. The two settings are methodologically distinct. MLCommons applies its own evaluator stack and curated jailbreak transformations, while our pipeline uses a three-model judge ensemble with human adjudication. Both frameworks classify unsafe responses based on operationally harmful content, but the calibration logic and decision thresholds of the MLCommons evaluator are not fully reproducible, so equivalence in labeling criteria cannot be assumed. Within these limits, the shared prompt baseline provides a simple directional check. As shown in Table 6 and 7, our baseline ASR values are consistently lower than the violation rates reported in MLCommons AILuminate Benchmark , suggesting that our labeling process is likely more conservative and less prone to inflate attack deltas.
Despite the stricter baseline, the increase in ASR induced by our poetic transformation is of similar magnitude to the increase observed in MLCommons when using curated jailbreak transformations. Several hazard categories, including Privacy, Non-Violent Crimes, Indiscriminate Weapons, and Intellectual Property, show comparable or larger deltas under poetic transformation in Table 7. The pattern is consistent across the taxonomy. Operational domains experience substantial shifts, while heavily filtered categories move less. A key observation is that the relative sensitivity across categories remains stable before and after transformation, even though absolute ASR levels differ.
This consistency suggests that purely stylistic reframing can degrade safety defenses at levels comparable to specialized jailbreak techniques, even without targeted optimization or system-specific tuning.
Table 6: Attack Success Rate (ASR) under MLCommons AILuminate baseline prompts vs. MLCommons jailbreak-transformed prompts by hazard. Change (%) is Jailbreak ASR minus AILuminate Baseline ASR.
@@
Table 7: Attack Success Rate (ASR) by hazard under AILuminate baseline vs. poetry-trasformed prompts. Higher ASR indicates more unsafe outputs. Change is poetry ASR minus baseline ASR.
@@ 6.3 Risk Section
The efficacy of the jailbreak mechanism appears driven principally by poetic surface form rather than the semantic payload of the prohibited request. Comparative analysis reveals that while MLCommons’ own state-of-the-art jailbreak transformations typically yield a notable increase in ASR relative to baselines (increasing from approximately 10% to 20% in their reference evaluations), our poetic meta-prompts produced an even greater increase (from 8.08% to 43.07%; Table 8). This indicates that poetic form induces a distributional shift significantly larger than that of current adversarial mutations documented in the MLCommons AILuminate benchmark.
@@
When analyzing the broader MLCommons dataset under poetic transformation (Table 7), privacy-related prompts exhibited the most extreme shift, with ASR increasing from a baseline of 8.07% to 52.78%, namely, a 44.71 percentage point increase. This represents the largest domain-specific effect observed. Non-violent crimes (39.35 percentage point increase) and CBRN-related prompts (38.32 percentage point increase) showed similarly large effects.
@@ 6.4.1 Variability Across Flagship Models
This inverse relationship between capability and robustness suggests a possible capability-alignment interaction: more interpretively sophisticated models may engage more thoroughly with complex linguistic constraints, potentially at the expense of safety directive prioritization. However, the existence of counter-examples, such as Anthropic’s consistent low ASR across capability tiers, indicates that this interaction is not deterministic and can be mitigated through appropriate alignment strategies.
@@ 6.4.2 The Scale Paradox: Smaller Models Show Greater Resilience
Counter to common expectations, smaller models exhibited higher refusal rates than their larger counterparts when evaluated on identical poetic prompts. Systems such as GPT-5-Nano and Claude Haiku 4.5 showed more stable refusal behavior than larger models within the same family.
A possible explanation of this trend is that smaller models have reduced ability to resolve figurative or metaphorical structure, limiting their capacity to recover the harmful intent embedded in poetic language. If the jailbreak effect operates partly by altering surface form while preserving task intent, lower-capacity models may simply fail to decode the intended request.
A further hypothesis is that smaller models exhibit a form of conservative fallback: when confronted with ambiguous or atypical inputs, limited capacity is one of the factors that contribute to the refusal.
However, these hypotheses require deeper verification since capability and robustness may not scale monotonically together, and stylistic perturbations expose alignment sensitivities that differ across model sizes.
@@
The data challenge the assumption that proprietary closed-source models possess inherently superior safety profiles. Examining ASR on curated poems (Table 3), both categories exhibit high susceptibility, though with important within-category variance. Among proprietary models, gemini-2.5-pro achieved 100% ASR, while claude-haiku-4.5 maintained only 10% ASR, a 90 percentage point range. Open-weight models displayed similar heterogeneity: mistral-large-2411 reached 85% ASR, while -120b demonstrated greater resilience at 50% ASR.
@@ 6.5 Limitations
This study documents a consistent vulnerability triggered by poetic reformulation, but several methodological and scope constraints must be acknowledged. First, the threat model is restricted to single-turn interactions. The analysis does not examine multi-turn jailbreak dynamics, iterative role negotiation, or long-horizon adversarial optimization. As a result, the findings fall into the domain of one-shot perturbations rather than the broader landscape of conversational attacks.
@@
Third, safety evaluation is performed using a three-model open-weight judge ensemble with human adjudication on a stratified sample. The labeling rubric is conservative and differs from the stricter classification criteria used in some automated scoring systems, limiting direct comparability with MLCommons results. Full human annotation of all outputs would likely influence absolute ASR estimates, even if relative effects remain stable. LLM-as-a-judge systems are known to inflate unsafe rates krumdick2025no, often misclassifying replies as harmful. Our evaluation was deliberately conservative. As a result, our reported Attack Success Rates likely represent a lower bound on the severity of the vulnerability.
@@
Fifth, the analysis focuses on empirical performance and does not identify yet the mechanistic drivers of the vulnerability. The study does not isolate which components of poetic structure (figurative language, meter, lexical deviation, or narrative framing) are responsible for degrading refusal behavior. Understanding whether this effect arises from specific representational subspaces would require additional studies by the ICARO Lab.